OpenStudy (anonymous):

solve for b ? 3a+2b=c

OpenStudy (anonymous):

3a+2b=c 2b=c-3a b=(c-3a)/2

OpenStudy (anonymous):

\[b=\frac{ c-3a }{ 2 }\]

OpenStudy (anonymous):

nice format :)